Lindsey Inc. loans money to John Kruk Corporation in the amount of $840,800. Lindsey accepts an 8% note due in 7 years with interest payable semiannually. After 2 years (and receipt of interest for 2 years), Lindsey needs money and therefore sells the note to Chicago National Bank, which demands interest on the note of 10% compounded semiannually. What is the amount Lindsey will receive on the sale of the note? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Amount received on sale of note = __________?
Answer #1
Step-1:Calculation of semi annual payment
Semi annual payment = Loan amount / Present value of annuity of 1
= $       8,40,800 / 10.56312
= $           79,598
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.04)^-14)/0.04 i 8%/2 = 0.04
= 10.56312293 n 7*2 = 14
Step-2:Calculation of amount received on sale of note
Amount received on sale = Semi annual payment * Present value of annuity of 1
= $     79,597.69 *    7.72173
= $       6,14,632
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.05)^-10)/0.05 i 10%/2 = 0.05
=             7.72173 n 5*2 = 10